# Modelling conductive thermal transport in three-dimensional fibrous media with fiber-to-fiber contacts

### v1

@@ III.2 Unified description of the evolution of the solid conductivity
As the conductivity of fibers $k_{fib}$ is fixed in the simulations discussed in the previous section, changing the value of the contact resistance $R_{k}$ amounts to changing the resistance ratio $R_{k}/R_{fib}$ where $R_{fib}=\frac{4l}{k_{fib}\pi d^{2}}$ is the intrinsic resistance of fibers. We verified by additional simulations that modifying the value of $R_{k}/R_{fib}$ by increasing $R_{k}$ or decreasing $k_{fib}$ gives the same results, for an identical network structure. Besides the resistance ratio, our results demonstrate that fiber parameters also participate in the control of conduction in the network. A relevant measure of this interplay between the intrinsic and contact resistances that takes into account the network geometry is given by a dimensionless parameter $r$ introduced in a theoretical model by Zhao et al. [48] and Volkov and Zhigilei [46, 49]:
@@
where $\langle N_{c}\rangle$ is the average number of contacts per fiber, $\langle H\rangle$ is the average vertical distance between two centers of fibers in contact, and $\langle\lvert\cos\theta\rvert\rangle$ is the average cosine of the polar angle of fibers. As highlighted by the authors, the theoretical models for 2D and 3D fibrous networks are similar, as a 3D network becomes 2D if the azimuthal angle reaches $\varphi=\pi/2$. Note that this transition from 3D to 2D networks is not investigated in this study as the azithumal angle is always chosen from a uniform distribution: the distribution of polar angles $\theta$ is the only one that can change. According to this model, the solid conductivity in a 3D fibrous network is given by:
$k_{cond,solid}=\frac{1}{4}k_{fib}\pi d^{2}\langle n_{z}\rangle\langle\lvert\cos\theta\rvert\rangle\frac{1}{1+r}=k_{0,th}\frac{1}{1+r}$ (4)
where $\langle n_{z}\rangle$ is the average areal number density of fibers through a horizontal section, and $k_{0,th}=\frac{1}{4}k_{fib}\pi d^{2}\langle n_{z}\rangle\langle\lvert\cos\theta\rvert\rangle$ is the theoretical value of the conductivity at $R_{k}=0$. This model is based on the assumptions of linearity of the temperature distributions along fibers, and of non-correlations between the temperatures of two fibers in contact.
@@
Figure 8: Solid thermal conductivity normalized to $k_{0,th}$, the theoretical conductivity at $R_{k}=0$, as a function of the characteristic dimensionless parameter $r$, for (a) different volume fractions $V_{f}$, (b) different aspect ratios $a$, and (c) different fiber orientations governed by the mean angle to heat flux direction $\bar{\theta}$. The base parameters are $V_{f}=0.01$, $a=100$ and $\bar{\theta}=57.3^{\circ}$. The black dashed line shows the predictions of Eq. (4).
compares the predictions of Eq. (4) to simulation data collected in this study, for different fiber volume fractions, aspect ratios, and orientations. Note that here, only networks with fiber mean angles between 57 and 86, i.e., with $\beta>1$, are considered, which is the only regime representative of real-world insulation materials. Fig. 8(a) indicates that the agreement between simulations and theory is improved when the volume fraction increases, as mentioned above. Similarly, Fig. 8(b) demonstrates a comparable trend concerning the aspect ratio. Note that this parameter has an effect independently of the individual values of $l$ and $d$ and is an important quantity in the conductivity scaling, which is consistent with our previous hypothesis. According to Fig. 8(c), fiber orientation does not have any effect on the relative variation of $k_{cond,solid}$ with $r$, indicating that the accordance between theory and simulation is fully determined by the values of $V_{f}$ and $a$. Overall, Fig. 8 shows that the above theoretical model seems to be relevant for 3D fibrous networks in the case of high volume fractions ($V_{f}>0.05$ when $a=100$) and high aspect ratios ($a>450$ when $V_{f}=0.01$), when adopting representative values of the properties of glasswool for the other parameters.
Fig. 9
@@
Figure 9: Solid thermal conductivity normalized to $k_{0}$, the simulated conductivity at $R_{k}=0$, as a function of the characteristic dimensionless parameter $r$, for different volume fractions $V_{f}$, aspect ratios $a$ and mean angles $\bar{\theta}$, the base parameters being $V_{f}=0.01$, $a=100$ and $\bar{\theta}=57.3^{\circ}$. The black dashed line shows the predictions of Eq. (4).
@@
represents, for different fiber parameters, the solid conductivity normalized to $k_{0}$, its value at $R_{k}=0$, as a function of the characteristic dimensionless parameter $r$. Simulated data is compared with the prediction $\frac{1}{1+r}$ given by Eq. (4). One can see that all markers corresponding to the simulation data collapse into a single master curve, regardless of the fiber volume fraction, aspect ratio or mean orientation, that is in good agreement with the theoretical model. This comparison validates, for any value of these network parameters, the consideration of $r$ as the characteristic parameter governing the evolution of solid conductivity as the contact resistance (or the intrinsic conductivity or fibers) changes. In the general case of a 3D fibrous assembly represented by straight lines, this quantity correctly describes the joint effect of resistance ratio and geometric properties. Obtaining this master curve also indicates that the deviation observed in Fig. 8(a) and (b) between model and simulation corresponds to a variable offset in the solid conductivity. Hence, a correction term appears to be needed in the theoretical prediction $k_{0,th}=\frac{1}{4}k_{fib}\pi d^{2}\langle n_{z}\rangle\langle\lvert\cos\theta\rvert\rangle$ to accurately describe fibrous media with low $a$ or $V_{f}$. Contrary to fiber orientation, these two parameters are found to have a significant effect on the average number of contacts per fiber $\langle N_{c}\rangle$, indicating the accordance with the model from Zhao et al. [48] and Volkov and Zhigilei [46] may be determined by the degree of connectivity of the simulated network. Indeed, when the aspect ratio or the density is low, the connectivity is low and the correlations between fiber temperatures may become non-negligible [47]. To validate this hypothesis, a modified version of the theoretical framework is derived in Appendix B, taking into account the contributions of correlations. This new model results in the following expression for the solid conductivity:
@@
$h=1+\frac{\langle\delta T_{ij}\rangle}{\nabla T_{z}\langle H\rangle}$ (6)
where $\langle\delta T_{ij}\rangle$ represents the average contribution of correlations in the temperature difference at junctions, and $\nabla T_{z}=\Delta T/L$ is the global temperature gradient imposed in the sample.
@@
Figure 10: Ratio of $k_{0}$, the simulated conductivity at $R_{k}=0$, to $k_{0,th}$, its corresponding theoretical prediction by the initial model according to Eq. (4), as a function of the average number of contacts per fiber $\langle N_{c}\rangle$. Colored symbols correspond to simulation results presented in Fig. 8(a) and 8(b), i.e., with different aspect ratios and volume fractions. The grey markers and the associated dashed curve represent the predictions of $h$ calculated from Eqs. (6) and (7).
. In this graph, the colored symbols show $k_{0}/k_{0,th}$, the ratio of the simulated conductivity at $R_{k}=0$ to its corresponding estimation by the initial model, as a function of $\langle N_{c}\rangle$. Note that this data corresponds to that of Fig. 8(a) and (b) obtained for different aspect ratios and volume fractions. It shows that the evolution of $k_{0}/k_{0,th}$ can be described by a single curve that depends only on $\langle N_{c}\rangle$, independently of the values of $a$ and $V_{f}$. This finding confirms that the observed reduction in the simulated conductivity compared with that predicted by the initial model is driven by network connectivity, accurately quantified by the average number of contacts per fiber. It can be qualitatively explained by the fact that in a highly-connected network, almost all fibers contribute to heat transfer, whereas only a small proportion of them do so in a poorly-connected one. Furthermore, grey symbols in Fig. 10 represent estimations of $h$ obtained by Eq. (6), for different aspect ratios and densities as well. Calculating $h$ only necessitates the additional measurement of $\langle\delta T_{ij}\rangle$ in simulations. This is achieved using the following expression extracted from the derivation of the modified theoretical model (see Appendix B):
@@
where $\langle|\Delta T_{ij}|\rangle$ is the average temperature difference between two fibers at a junction, and $\langle\frac{dT}{dz}\rangle$ is the average temperature gradient in fibers along the $z$-direction. The quantity $h$ should correspond to the corrective prefactor $k_{0}/k_{0,th}$ needed to account for low-connectivity networks. Fig. 10 shows good agreement between these two quantities, despite notable fluctuations in the calculated values of $h$. Note that finding $0\leq h\leq 1$ is a direct result of $\langle\delta T_{ij}\rangle$ being a negative value, indicating temperature differences at junctions are lower in average in the modified model. Indeed, considering the correlations results in a reduction of temperature differences between closely positioned fibers, especially those in direct contact. A deviation is observed at high $\langle N_{c}\rangle$, with $k_{0}/k_{0,th}$ reaching values higher than 1. As the connectivity increases, the prefactor of the theoretical model should tend towards 1: this suggests that $k_{0}/k_{0,th}$ may be slightly overestimated by our simulations, while the estimate for $h$ appears to have rectified this discrepancy, indicating a higher level of accuracy. The dashed line plotted in Fig. 10 represents the best fit of the numerical data obtained for $h$ by a rational function $h=C_{1}\frac{\langle N_{c}\rangle-C_{2}}{\langle N_{c}\rangle-C_{3}}$ where $C_{1}$, $C_{2}$ and $C_{3}$ are adjustable parameters; the choice of such a function being motivated by the form of Eq. (6). This best fit gives $C_{1}\approx C_{3}\approx 1$ and $C_{2}\approx 2.18$, confirming our data is consistent with $h\rightarrow 1$ when $\langle N_{c}\rangle\rightarrow\infty$. The obtained best fit function can be rewritten as:
$h=1-\frac{\langle N_{c}\rangle_{l}-1}{\langle N_{c}\rangle-1}$ (8)
@@
where $\langle N_{c}\rangle_{l}=C_{2}=2.18$ is the minimum number of contacts per fiber ensuring a conductive network, i.e., that corresponds to the percolation threshold. Note that as the minimum number of contacts for each fiber is 2 to form a fully-connected network where thermal transfers can be solved, it was expected that the minimum value averaged at the network scale would indeed be slightly higher. In previous studies [60, 46], the average number of contacts per fiber has been expressed as a function of the input parameters: for 3D networks, $\langle N_{c}\rangle=2aV_{f}$. Therefore, Fig. 10 can be interpreted as demonstrating a percolative behavior for 3D networks governed by both volume fraction and aspect ratio: in particular, when $a$ approaches 0, regardless of the value of $V_{f}$, the network conductivity becomes zero, as it falls below the percolation threshold. As stated in Ref. [49], this constitutes a major difference with 2D networks, for which the conductivity can remain finite for very small aspect ratios: this is due to the topological difference between 2D and 3D networks.
@@
$\langle N_{c}\rangle-1$ can be interpreted as the average number of conducting segments per fiber, each one of them being delimited by contact points. Therefore, the deviation from unity of $h$, which is due to the correlation term $\frac{\langle\delta T_{ij}\rangle}{\nabla T_{z}\langle H\rangle}$ according to Eq. (6), can be entirely explained by this number of segments per fiber. As it decreases and approaches its percolation limit value, temperature distributions along fibers exhibit more pronounced piecewise-linear behavior, therefore increasingly deviating from a linear gradient, as it can be seen in Fig. S4 (Supplementary Material) [50]. This explains the growing importance of the impact of correlations in the approximate decomposition of the temperature of a fiber made in the theoretical model (see Eq. (19) in Appendix B). More generally, this result shows that the solid conductivity in any 3D network of straight fibers can be predicted based on geometric considerations only, using Eqs. (5) and (8), which do not require the resolution of the heat transfer problem.
@@
Eq. (8) also allows to compute a threshold value for $\langle N_{c}\rangle$, above which the initial model from Zhao et al. [48] and Volkov and Zhigilei [46] can be considered as sufficient. For a 5% tolerance, this threshold is $\langle N_{c}\rangle_{t,5\%}=24.6$. This value may be seen as a separation between two characteristic connectivity regimes for a 3D fibrous assembly, generalizing the low-density and high-density regimes described in section III.1.1, where we gave a rough estimate for a given case with fixed fiber dimensions. In addition, the expression introduced previously linking the average number of contacts per fiber to the network input parameters in 3D networks, $\langle N_{c}\rangle=2aV_{f}$, can be used to estimate before any numerical generation the value of $h$ for a given set of network parameters. For an average aspect ratio $a=100$, representative of fibers in glasswool, the threshold $\langle N_{c}\rangle_{t,5\%}=24.6$ converts into a minimum volume fraction $V_{f_{t,5\%}}=0.123$. This result indicates that to accurately represent most of glasswool insulation mats, that exhibit lower densities, the model taking into account correlations that we derived and validated in this study should be considered, underlining the interest of this work.
@@
In practice, to predict the thermal conductivity of a specific fibrous insulation material using this model, according to Eqs. (5) and (8), the input parameters $d$, $l$, $\langle\lvert\cos\theta\rvert\rangle$, $V_{f}$, $R_{k}$ and $k_{f}$ must be known. The three additional quantities $\langle N_{c}\rangle$, $\langle H\rangle$ and $\langle n_{z}\rangle$ necessary to calculate $k_{cond,solid}$ can then be obtained from the numerical generation of a fibrous network characterized by the above parameters, as it is done in this study. For example, Fig. 11
@@
Figure 11: Simulated (plain line) and theoretically estimated (dashed line) solid conductivity as a function of estimated glasswool density, calculated based on fiber volume fraction and considering the additional presence of binder in real samples. The constant input parameters are $l=1$ mm, $d=10~{}\mu$m and $\bar{\theta}=79.9^{\circ}$, with fiber volume fractions $V_{f}$ ranging from to 0.006 to 0.03. The solid conductivities estimated using the theoretical model presented above are obtained using Eqs. (5) and (8).

### v2

@@ III.2 Unified description of the evolution of the solid conductivity
As the conductivity of fibers $k_{fib}$ is fixed in the simulations discussed in the previous section, changing the value of the contact resistance $R_{k}$ amounts to changing the resistance ratio $R_{k}/R_{fib}$ where $R_{fib}=\frac{4l}{k_{fib}\pi d^{2}}$ is the intrinsic resistance of fibers. We verified with additional simulations that modifying the value of $R_{k}/R_{fib}$ by increasing $R_{k}$ or decreasing $k_{fib}$ gives the same results, for an identical network structure. Besides the resistance ratio, our results demonstrate that fiber parameters also participate in the control of conduction in the network. A relevant measure of this interplay between the intrinsic and contact resistances that takes into account the network geometry is given by a dimensionless parameter $r$ introduced in a theoretical model by Zhao et al. [48] and Volkov and Zhigilei [46, 49]:
@@
. Here $\langle N_{c}\rangle$ is the average number of contacts per fiber, $\langle H\rangle$ is the average vertical distance between two centers of fibers in contact, and $\langle\lvert\cos\theta\rvert\rangle$ is the average cosine of the polar angle of fibers. As highlighted by the authors, the theoretical models for 2D and 3D fibrous networks are similar, as a 3D network becomes 2D if the azimuthal angle reaches $\varphi=\pi/2$. Note that this transition from 3D to 2D networks is not investigated in this study as the azimuthal angle is always chosen from a uniform distribution: the distribution of polar angles $\theta$ is the only one that can change. According to this model, the solid conductivity in a 3D fibrous network is given by
$k_{cond,solid}=\frac{1}{4}k_{fib}\pi d^{2}\langle n_{z}\rangle\langle\lvert\cos\theta\rvert\rangle\frac{1}{1+r}=k_{0,th}\frac{1}{1+r},$ (4)
where $\langle n_{z}\rangle$ is the average areal number density of fibers through a horizontal section, and $k_{0,th}=\frac{1}{4}k_{fib}\pi d^{2}\langle n_{z}\rangle\langle\lvert\cos\theta\rvert\rangle$ is the theoretical value of the conductivity at $R_{k}=0$. This model is based on the assumptions of linearity of the temperature distributions along fibers, and of noncorrelations between the temperatures of two fibers in contact.
Figure 8
Figure 8: Solid thermal conductivity normalized to $k_{0,th}$, the theoretical conductivity at $R_{k}=0$, as a function of the characteristic dimensionless parameter $r$ for (a) different volume fractions $V_{f}$, (b) different aspect ratios $a$, and (c) different fiber orientations governed by the mean angle to heat flux direction $\bar{\theta}$. The base parameters are $V_{f}=0.01$, $a=100$, and $\bar{\theta}=57.3^{\circ}$. The black dashed line shows the predictions of Eq. (4).
compares the predictions of Eq. (4) to simulation data collected in this study for different fiber volume fractions, aspect ratios, and orientations. Note that here only networks with fiber mean angles between 57 and 86, i.e., with $\beta>1$, are considered, which is the only regime representative of real-world insulation materials. Figure 8(a) indicates that the agreement between simulations and theory is improved when the volume fraction increases, as mentioned above. Similarly, Fig. 8(b) demonstrates a comparable trend concerning the aspect ratio. Note that this parameter has an effect independently of the individual values of $l$ and $d$ and is an important quantity in the conductivity scaling, which is consistent with our previous hypothesis. According to Fig. 8(c), fiber orientation does not have any effect on the relative variation of $k_{cond,solid}$ with $r$, indicating that the accordance between theory and simulation is fully determined by the values of $V_{f}$ and $a$. Overall, Fig. 8 shows that the above theoretical model seems to be relevant for 3D fibrous networks in the case of high volume fractions ($V_{f}>0.05$ when $a=100$) and high aspect ratios ($a>450$ when $V_{f}=0.01$), when adopting representative values of the properties of glasswool for the other parameters.
@@
Figure 9: Solid thermal conductivity normalized to $k_{0}$, the simulated conductivity at $R_{k}=0$, as a function of the characteristic dimensionless parameter $r$ for different volume fractions $V_{f}$, aspect ratios $a$, and mean angles $\bar{\theta}$. The base parameters are $V_{f}=0.01$, $a=100$, and $\bar{\theta}=57.3^{\circ}$. The black dashed line shows the predictions of Eq. (4).
@@
represents, for different fiber parameters, the solid conductivity normalized to $k_{0}$, its value at $R_{k}=0$, as a function of the characteristic dimensionless parameter $r$. Simulated data are compared with the prediction $\frac{1}{1+r}$ given by Eq. (4). One can see that all markers corresponding to the simulation data collapse into a single master curve, regardless of the fiber volume fraction, aspect ratio, or mean orientation, which is in good agreement with the theoretical model. This comparison validates, for any value of these network parameters, the consideration of $r$ as the characteristic parameter governing the evolution of the solid conductivity as the contact resistance (or the intrinsic conductivity of fibers) changes. In the general case of a 3D fibrous assembly represented by straight lines, this quantity correctly describes the joint effect of resistance ratio and geometric properties. Obtaining this master curve also indicates that the deviation observed in Figs. 8(a) and (b) between the model and simulation corresponds to a variable offset in the solid conductivity. Hence, a correction term appears to be needed in the theoretical prediction $k_{0,th}=\frac{1}{4}k_{fib}\pi d^{2}\langle n_{z}\rangle\langle\lvert\cos\theta\rvert\rangle$ to accurately describe fibrous media with low $a$ or $V_{f}$. Contrary to fiber orientation, these two parameters are found to have a significant effect on the average number of contacts per fiber $\langle N_{c}\rangle$, indicating that the accordance with the model from Zhao et al. [48] and Volkov and Zhigilei [46] may be determined by the degree of connectivity of the simulated network. Indeed, when the aspect ratio or the density is low, the connectivity is low and the correlations between fiber temperatures may become non-negligible [47]. To validate this hypothesis, a modified version of the theoretical framework is derived in Appendix B, taking into account the contributions of correlations. This new model results in the following expression for the solid conductivity:
@@
$h=1+\frac{\langle\delta T_{ij}\rangle}{\nabla T_{z}\langle H\rangle}.$ (6)
Here $\langle\delta T_{ij}\rangle$ represents the average contribution of correlations in the temperature difference at junctions, and $\nabla T_{z}=\Delta T/L$ is the global temperature gradient imposed in the sample.
@@
Figure 10: Ratio of $k_{0}$, the simulated conductivity at $R_{k}=0$, to $k_{0,th}$, its corresponding theoretical prediction by the initial model according to Eq. (4), as a function of the average number of contacts per fiber $\langle N_{c}\rangle$. Colored symbols correspond to simulation results presented in Figs. 8(a) and 8(b), i.e., with different aspect ratios and volume fractions. The gray markers and the associated dashed curve represent the predictions of $h$ calculated from Eqs. (6) and (7).
. In this graph, the colored symbols show $k_{0}/k_{0,th}$, the ratio of the simulated conductivity at $R_{k}=0$ to its corresponding estimation by the initial model, as a function of $\langle N_{c}\rangle$. Note that these data correspond to those of Figs. 8(a) and (b) obtained for different aspect ratios and volume fractions. They show that the evolution of $k_{0}/k_{0,th}$ can be described by a single curve that depends only on $\langle N_{c}\rangle$, independently of the values of $a$ and $V_{f}$. This finding confirms that the observed reduction in the simulated conductivity compared with that predicted by the initial model is driven by network connectivity, accurately quantified by the average number of contacts per fiber. It can be qualitatively explained by the fact that in a highly connected network, almost all fibers contribute to heat transfer, whereas only a small proportion of them do so in a poorly connected one. Furthermore, the gray symbols in Fig. 10 represent estimations of $h$ obtained by Eq. (6), for different aspect ratios and densities as well. Calculating $h$ only necessitates the additional measurement of $\langle\delta T_{ij}\rangle$ in simulations. This is achieved using the following expression extracted from the derivation of the modified theoretical model (see Appendix B):
@@
with $\langle|\Delta T_{ij}|\rangle$ the average temperature difference between two fibers at a junction, and $\langle\frac{dT}{dz}\rangle$ the average temperature gradient in fibers along the $z$ direction. The quantity $h$ should correspond to the corrective prefactor $k_{0}/k_{0,th}$ needed to account for low-connectivity networks. Figure 10 shows good agreement between these two quantities, despite notable fluctuations in the calculated values of $h$. Note that finding $0\leq h\leq 1$ is a direct result of $\langle\delta T_{ij}\rangle$ being a negative value, indicating that temperature differences at junctions are lower in average in the modified model. Indeed, considering the correlations results in a reduction of temperature differences between closely positioned fibers, especially those in direct contact. A deviation is observed at high $\langle N_{c}\rangle$, with $k_{0}/k_{0,th}$ reaching values higher than 1. As the connectivity increases, the prefactor of the theoretical model should tend towards 1: this suggests that $k_{0}/k_{0,th}$ may be slightly overestimated by our simulations, while the estimate for $h$ appears to have rectified this discrepancy, indicating a higher level of accuracy. The dashed line plotted in Fig. 10 represents the best fit of the numerical data obtained for $h$ by a rational function $h=C_{1}\frac{\langle N_{c}\rangle-C_{2}}{\langle N_{c}\rangle-C_{3}}$ where $C_{1}$, $C_{2}$, and $C_{3}$ are adjustable parameters; the choice of such a function is motivated by the form of Eq. (6). This best fit gives $C_{1}\approx C_{3}\approx 1$ and $C_{2}\approx 2.18$, confirming that our data are consistent with $h\rightarrow 1$ when $\langle N_{c}\rangle\rightarrow\infty$. The obtained best-fit function can be rewritten as
$h=1-\frac{\langle N_{c}\rangle_{l}-1}{\langle N_{c}\rangle-1},$ (8)
@@
where $\langle N_{c}\rangle_{l}=C_{2}=2.18$ is the minimum number of contacts per fiber ensuring a conductive network, i.e., that corresponds to the percolation threshold. Note that as the minimum number of contacts for each fiber is 2 to form a fully connected network where thermal transfers can be solved, it was expected that the minimum value averaged at the network scale would indeed be slightly higher. In previous studies [60, 46], the average number of contacts per fiber has been expressed as a function of the input parameters: for 3D networks, $\langle N_{c}\rangle=2aV_{f}$. Therefore, Fig. 10 can be interpreted as demonstrating percolative behavior for 3D networks governed by the both volume fraction and aspect ratio; in particular, when $a$ approaches 0, regardless of the value of $V_{f}$, the network conductivity becomes zero, as it falls below the percolation threshold. As stated in Ref. [49], this constitutes a major difference with 2D networks, for which the conductivity can remain finite for very small aspect ratios: this is due to the topological difference between 2D and 3D networks.
@@
We can interpret $\langle N_{c}\rangle-1$ as the average number of conducting segments per fiber, each one of them being delimited by contact points. Therefore, the deviation from unity of $h$, which is due to the correlation term $\frac{\langle\delta T_{ij}\rangle}{\nabla T_{z}\langle H\rangle}$ according to Eq. (6), can be entirely explained by this number of segments per fiber. As it decreases and approaches its percolation limit value, temperature distributions along fibers exhibit more pronounced piecewise-linear behavior, therefore increasingly deviating from a linear gradient, as can be seen in Fig. S4 with the Supplemental Material [50]. This explains the growing importance of the impact of correlations in the approximate decomposition of the temperature of a fiber made in the theoretical model (see Eq. (19) in Appendix B). More generally, this result shows that the solid conductivity in any 3D network of straight fibers can be predicted based on geometric considerations only, using Eqs. (5) and (8), which do not require resolution of the heat transfer problem.
@@
Equation. (8) also allows us to compute a threshold value for $\langle N_{c}\rangle$, above which the initial model from Zhao et al. [48] and Volkov and Zhigilei [46] can be considered as sufficient. For a 5% tolerance, this threshold is $\langle N_{c}\rangle_{t,5\%}=24.6$. This value may be seen as a separation between two characteristic connectivity regimes for a 3D fibrous assembly, generalizing the low-density and high-density regimes described in Sec. III.1.1, where we gave a rough estimate for a given case with fixed fiber dimensions. In addition, the expression introduced previously linking the average number of contacts per fiber to the network input parameters in 3D networks, $\langle N_{c}\rangle=2aV_{f}$, can be used before any numerical generation to estimate the value of $h$ for a given set of network parameters. For an average aspect ratio $a=100$, representative of fibers in glasswool, the threshold $\langle N_{c}\rangle_{t,5\%}=24.6$ converts into a minimum volume fraction $V_{f_{t,5\%}}=0.123$. This result indicates that to accurately represent most glasswool insulation mats that exhibit lower densities, the model taking into account correlations that we derived and validated in this study should be considered, underlining the interest of this work.
@@
In practice, to predict the thermal conductivity of a specific fibrous insulation material using this model, according to Eqs. (5) and (8), the input parameters $d$, $l$, $\langle\lvert\cos\theta\rvert\rangle$, $V_{f}$, $R_{k}$, and $k_{f}$ must be known. The three additional quantities $\langle N_{c}\rangle$, $\langle H\rangle$, and $\langle n_{z}\rangle$ necessary to calculate $k_{cond,solid}$ can then be obtained from the numerical generation of a fibrous network characterized by the above parameters, as done in this study. For example, Fig. 11
@@
Figure 11: Simulated (plain line) and theoretically estimated (dashed line) solid conductivities as a function of the estimated glasswool density, calculated based on the fiber volume fraction and considering the additional presence of a binder in real samples. The constant input parameters are $l=1$ mm, $d=10~{}\mu$m, and $\bar{\theta}=79.9^{\circ}$, with fiber volume fractions $V_{f}$ ranging from 0.006 to 0.03. The solid conductivities estimated using the theoretical model presented above are obtained using Eqs. (5) and (8).